# references.bib

### v4

@@ 1 Introduction
Identifiability is a property which a model must satisfy for precise parameter inference, and identifiability analysis refers to a group of methods used to determine how well the parameters of a mathematical model can be estimated given the observed data [Cobelli1980], which may vary widely in quality and quantity across different practical applications. Methods of identifiability analysis are typically classified in terms of whether they deal with structural or practical identifiability [Hines2014]. Structural identifiability focuses on the question of whether different parameter values generate different probability distributions of the observable variables [Meshkat2009, Meshkat2015]. The implication of this is that, with access to an infinite amount of ideal, noise-free data, it would then be possible to precisely estimate the model parameters. As such, structural identifiability is solely concerned with analyzing the structure of the mathematical model. Structural identifiability is often assessed using software that typically use Lie derivatives to generate a system of input–output equations, and the solvability properties of this system provide information about structural identifiability [Chis2011, Díaz-Seoane2022, Ligon2018]. In contrast, practical identifiability analysis involves working with noisy, incomplete data, and exploring the extent to which parameter values can be confidently estimated, given these data. In particular, it usually entails fitting a mathematical model to data and then exploring the extent to which the fit of the model to the data changes as the parameters are varied. Such practical identifiability analysis can be performed either locally near a given point, such as near the parameter values that provide the best model fit, or globally over the extended parameter space. A common tool for assessing practical identifiability is the profile likelihood [Frohlich2014, Raue2009], which is the main focus of this article.
@@ 3 Modeling with PDEs
where $D>0$ [L2/T] is the diffusion coefficient, and $v$ [L/T] is the advective velocity. We have not specified the units of $D$ or $v$, instead we use standard notation [Barenblatt2003] to indicate that the diffusion coefficient has dimensions of length squared per time, and the advective velocity has dimensions of length per time. This mathematical model is widely used in various physical, chemical and biological applications, such as the study of the dispersion of dissolved solutes (e.g. nutrients, pollutants) in porous media [Zheng2002]. We will consider the solution of Equation (5) for the initial condition $u(x,0)=u_{b}+u_{0}$ for $|x|<h$, and $u(x,0)=u_{b}$ for $|x|>h$, where $h>0$ is the half-width of the initial condition about the origin. We interpret $u_{b}\geq 0$ as a uniform background concentration of $u$, and we are particularly interested in the spatial spreading of $u$ that arises when an additional amount of $u$ is placed uniformly within distance $h>0$ from the origin, giving $u(x,0)=u_{b}+u_{0}$ for $|x|<h$, where $u_{0}>0$. This initial condition with $h=50$ is plotted in Figure 4(a) where we have a background concentration $u_{b}=1$ in the far-field, and locally within the interval $-50<x<50$ we have $u(x,0)=2$ (i.e. $u_{0}=1$). The solution of the mathematical model describes how the initial density profile evolves as a result of combined advection and diffusive transport as a function of position $x$ and time $t$.
Figure 4: (a) Initial condition $u(x,0)$ (solid green) with $u_{b}=u_{0}=1$ and $h=50$ superimposed with the PDE solution at $t=50$, $u(x,50)$ (solid red) with $D=10$ and $v=1$. (b) Synthetic data (blue dots) showing observations $u^{\rm{obs}}(x,50)$ at $x=-200,-195,-190,\ldots,195,200$ superimposed with the MLE solution (solid red) with $\hat{\boldsymbol{\theta}}=(\hat{u_{0}},\hat{h},\hat{D},\hat{v})^{\mathsf{T}}=(0.914,53.693,8.671,1.025)^{\mathsf{T}}$ and the 95% prediction interval (green shaded region). The data are obtained by evaluating Equation (5) with $(u_{0},h,D,v)^{\mathsf{T}}=(1,50,10,1)^{\mathsf{T}}$ at $x=-200,-195,-190,\ldots,195,200$ and corrupting each data point with additive Gaussian noise with $\sigma=0.05$. (c)–(f) Profile likelihood functions for $u_{0}$, $h$, $D$ and $v$, as indicated (solid red). Each profile indicates the MLE (solid blue) and the $95\%$ threshold ¯$\bar{\ell}^{*}=-\Delta_{0.95,1}/2=-1.921$ (solid gold). The 95% confidence intervals are $u_{0}\in[0.839,1.156]$, $h\in[42.939,58.043]$, $D\in[6.227,13.686]$ and $v\in[0.990,1.060]$.
@@
$u(x,t)=u_{b}+\dfrac{u_{0}}{2}\left[\textrm{erf}\left(\dfrac{h-(x-vt)}{2\sqrt{Dt}}\right)+\textrm{erf}\left(\dfrac{h+(x-vt)}{2\sqrt{Dt}}\right)\right],$ (6)
@@
This PDE example highlights an important shortcoming of working with the additive Gaussian noise model. A mathematical property of the exact solution, Equation (6), is that $u_{b}<u(x,t)<u_{0}+u_{b}$ for $t>0$ and $-\infty<x<\infty$ since the parabolic PDE model obeys a maximum principal [Protter1967]. Our data in Figure 4(b) clearly violates this property as we have $u^{\rm{obs}}(x,t)<u_{b}$ at several locations. This issue is even more concerning if we consider the realistic case of having no background concentration by setting $u_{b}=0$. Under these circumstances working with additive Gaussian noise is clearly unsatisfactory since this leads to $u^{\rm{obs}}(x,t)<0$ which is physically impossible when $u$ represents a concentration that is non-negative by definition.
One way to address this shortcoming is to work with a different noise model. In this case we could instead introduce a multiplicative noise model where the fluctuations are no longer constant or symmetric about the mean. For example, we could specify a noise model where the magnitude of the fluctuations is proportional to the mean of the distribution, $u^{\rm{obs}}(x_{i},t)\mid\boldsymbol{\theta}=(u(x_{i},t)\mid\boldsymbol{\theta})\eta_{i}$ where $\eta_{i}~{}\sim~{}\textrm{log-normal}(0,\sigma^{2})$ [Murphy2024]. Our previous examples with additive Gaussian noise models have a constant variability whereas the multiplicative log-normal noise model has variability that increases with $u^{\rm{obs}}(x,t)$. The log-normal noise model also has the attractive property that the variability vanishes as $u^{\rm{obs}}(x,t)~{}\to~{}0^{+}$. Figure 5(a) shows a solution of Equation (5) with $u_{b}=0$ and $(u_{0},h,D,v)^{\mathsf{T}}=(1,50,10,1)^{\mathsf{T}}$ at $t=50$, where the solution at $x=-200,-195,-190,\ldots,200$ is corrupted with multiplicative log-normal noise with $\sigma=0.2$. Here we see that the data is noise-free as $u\to 0^{+}$ and the variability is largest near $x=50$ where $u(x,t)$ is a maximum at this time.

### v5

@@ 1 Introduction
Identifiability is a property which a model must satisfy for precise parameter inference, and identifiability analysis refers to a group of methods used to determine how well the parameters of a mathematical model can be estimated given the observed data [Cobelli1980], which may vary widely in quality and quantity across different practical applications. Methods of identifiability analysis are typically classified in terms of whether they deal with structural or practical identifiability [Hines2014]. Structural identifiability focuses on the question of whether different parameter values generate different probability distributions of the observable variables [Meshkat2009, Meshkat2015, Simpson2024b]. The implication of this is that, with access to an infinite amount of ideal, noise-free data, it would then be possible to precisely estimate the model parameters. As such, structural identifiability is solely concerned with analyzing the structure of the mathematical model. Structural identifiability is often assessed using software that typically use Lie derivatives to generate a system of input–output equations, and the solvability properties of this system provide information about structural identifiability [Chis2011, Díaz-Seoane2022, Ligon2018]. In contrast, practical identifiability analysis involves working with noisy, incomplete data, and exploring the extent to which parameter values can be confidently estimated, given these data. In particular, it usually entails fitting a mathematical model to data and then exploring the extent to which the fit of the model to the data changes as the parameters are varied. Such practical identifiability analysis can be performed either locally near a given point, such as near the parameter values that provide the best model fit, or globally over the extended parameter space. A common tool for assessing practical identifiability is the profile likelihood [Frohlich2014, Raue2009, Simpson2024b], which is the main focus of this article.
@@ 3 Modeling with PDEs
where $D>0$ [L2/T] is the diffusion coefficient, and $v$ [L/T] is the advective velocity. We have not specified the units of $D$ or $v$, instead we use standard notation [Barenblatt2003] to indicate that the diffusion coefficient has dimensions of length squared per time, and the advective velocity has dimensions of length per time. This mathematical model is widely used in various physical, chemical and biological applications, such as the study of the dispersion of dissolved solutes (e.g. nutrients, pollutants) in porous media [Zheng2002]. We will consider the solution of Equation (5) for the initial condition $u(x,0)=u_{\textrm{b}}+u_{0}$ for $|x|<h$, and $u(x,0)=u_{\textrm{b}}$ for $|x|>h$, where $h>0$ is the half-width of the initial condition about the origin. We interpret $u_{\textrm{b}}\geq 0$ as a uniform background concentration of $u$, and we are particularly interested in the spatial spreading of $u$ that arises when an additional amount of $u$ is placed uniformly within distance $h>0$ from the origin, giving $u(x,0)=u_{\textrm{b}}+u_{0}$ for $|x|<h$, where $u_{0}>0$. This initial condition with $h=50$ is plotted in Figure 4(a) where we have a background concentration $u_{\textrm{b}}=1$ in the far-field, and locally within the interval $-50<x<50$ we have $u(x,0)=2$ (i.e. $u_{0}=1$). The solution of the mathematical model describes how the initial density profile evolves as a result of combined advection and diffusive transport as a function of position $x$ and time $t$.
Figure 4: (a) Initial condition $u(x,0)$ (solid green) with $u_{\textrm{b}}=u_{0}=1$ and $h=50$ superimposed with the PDE solution at $t=50$, $u(x,50)$ (solid red) with $D=10$ and $v=1$. (b) Synthetic data (blue dots) showing observations $u^{\rm{obs}}(x,50)$ at $x=-200,-195,-190,\ldots,195,200$ superimposed with the MLE solution (solid red) with $\hat{\boldsymbol{\theta}}=(\hat{u_{0}},\hat{h},\hat{D},\hat{v})^{\mathsf{T}}=(0.914,53.693,8.671,1.025)^{\mathsf{T}}$ and the 95% prediction interval (green shaded region). The data are obtained by evaluating Equation (5) with $(u_{0},h,D,v)^{\mathsf{T}}=(1,50,10,1)^{\mathsf{T}}$ at $x=-200,-195,-190,\ldots,195,200$ and corrupting each data point with additive Gaussian noise with $\sigma=0.05$. (c)–(f) Profile likelihood functions for $u_{0}$, $h$, $D$ and $v$, as indicated (solid red). Each profile indicates the MLE (solid blue) and the $95\%$ threshold ¯$\bar{\ell}^{*}=-\Delta_{0.95,1}/2=-1.921$ (solid gold). The 95% confidence intervals are $u_{0}\in[0.839,1.156]$, $h\in[42.939,58.043]$, $D\in[6.227,13.686]$ and $v\in[0.990,1.060]$.
@@
$u(x,t)=u_{\textrm{b}}+\dfrac{u_{0}}{2}\left[\textrm{erf}\left(\dfrac{h-(x-vt)}{2\sqrt{Dt}}\right)+\textrm{erf}\left(\dfrac{h+(x-vt)}{2\sqrt{Dt}}\right)\right],$ (6)
@@
This PDE example highlights an important shortcoming of working with the additive Gaussian noise model. A mathematical property of the exact solution, Equation (6), is that $u_{\textrm{b}}<u(x,t)<u_{0}+u_{\textrm{b}}$ for $t>0$ and $-\infty<x<\infty$ since the parabolic PDE model obeys a maximum principal [Protter1967]. Our data in Figure 4(b) clearly violates this property as we have $u^{\rm{obs}}(x,t)<u_{\textrm{b}}$ at several locations. This issue is even more concerning if we consider the realistic case of having no background concentration by setting $u_{\textrm{b}}=0$. Under these circumstances working with additive Gaussian noise is clearly unsatisfactory since this leads to $u^{\rm{obs}}(x,t)<0$ which is physically impossible when $u$ represents a concentration that is non-negative by definition.
One way to address this shortcoming is to work with a different noise model. In this case we could instead introduce a multiplicative noise model where the fluctuations are no longer constant or symmetric about the mean. For example, we could specify a noise model where the magnitude of the fluctuations is proportional to the mean of the distribution, $u^{\rm{obs}}(x_{i},t)\mid\boldsymbol{\theta}=(u(x_{i},t)\mid\boldsymbol{\theta})\eta_{i}$ where $\eta_{i}~{}\sim~{}\textrm{log-normal}(0,\sigma^{2})$ [Murphy2024]. Our previous examples with additive Gaussian noise models have a constant variability whereas the multiplicative log-normal noise model has variability that increases with $u^{\rm{obs}}(x,t)$. The log-normal noise model also has the attractive property that the variability vanishes as $u^{\rm{obs}}(x,t)~{}\to~{}0^{+}$. Figure 5(a) shows a solution of Equation (5) with $u_{\textrm{b}}=0$ and $(u_{0},h,D,v)^{\mathsf{T}}=(1,50,10,1)^{\mathsf{T}}$ at $t=50$, where the solution at $x=-200,-195,-190,\ldots,200$ is corrupted with multiplicative log-normal noise with $\sigma=0.2$. Here we see that the data is noise-free as $u\to 0^{+}$ and the variability is largest near $x=50$ where $u(x,t)$ is a maximum at this time.